In a large population 63% of the people have been vaccinated if 5 people are randomly selected what is the probability that at least one of them has been vaccinated


1
Expert's answer
2022-03-15T17:29:23-0400

This is a binomial distribution with n=5, p=0.63.

"P(X\ge1)=1-P(X=0)=1-C_5^00.63^0(1-0.63)^5=0.9931."